OpenStudy (anonymous):

what is the limit of (3-h)^-1 - 3^-1 / h as h approaches 0

OpenStudy (anonymous):

\[\lim \frac{ (3-h)^{-1} - 3^{-1}}{ h }\]

OpenStudy (anonymous):

as h approaches 0

OpenStudy (anonymous):

is th answer: indeterminate?

OpenStudy (anonymous):

No, you just need to manipulate it a bit: \[\begin{align*}\lim_{h\to0}\frac{\dfrac{1}{3-h}-\dfrac{1}{3}}{h}&=\lim_{h\to0}\frac{\dfrac{3}{3(3-h)}-\dfrac{3-h}{3(3-h)}}{h}\\ &=\lim_{h\to0}\frac{1}{h}\left(\frac{3-(3-h)}{3(3-h)}\right)\\ \end{align*}\]

OpenStudy (anonymous):

so the limit is 1

OpenStudy (anonymous):

That's not what I'm getting, no.

OpenStudy (anonymous):

you would canel out the (3-h)

OpenStudy (anonymous):

*cancel

OpenStudy (anonymous):

No, you wouldn't, that's not how dividing works:\[\lim_{h\to0}\frac{1}{h}\left(\frac{3-(3-h)}{3(3-h)}\right)=\lim_{h\to0}\frac{1}{h}\left(\frac{h}{3(3-h)}\right)=\lim_{h\to0}\frac{1}{3(3-h)}\]

OpenStudy (anonymous):

kk. that makes sense. Im stupid

OpenStudy (anonymous):

Don't worry about it, just takes some practice :)

OpenStudy (anonymous):

isnt there supposed to be a (-) in front of the h in the numerator?

OpenStudy (anonymous):

oh wait, is it 3-3+h?

OpenStudy (anonymous):

It's \(3-(3-h)=3-3+h=h\).

OpenStudy (anonymous):

so the answer is 1/9

OpenStudy (anonymous):

Yes, that's right

OpenStudy (anonymous):

thank you so much.